# Diffusion Lens: Interpreting Text Encoders in Text-to-Image Pipelines

### v1

@@ 1 Introduction
The text-to-image (T2I) diffusion pipeline comprises two main elements: the text encoder and the diffusion model. The text encoder converts a textual prompt into a latent representation, while the diffusion model utilizes this representation to generate the corresponding image. Several recent studies have delved into the internal workings of the diffusion model and the cross-attention mechanism that connects the two components Tang et al. (2023); Hertz et al. (2023); Orgad et al. (2023); Chefer et al. (2023a). Yet, while the text encoder is a key component of the pipeline with a large effect on image quality and text–image alignment Saharia et al. (2022), its internal mechanisms remain unexplored. Moreover, while there is a wide range of work that has analyzed general language model internals Belinkov and Glass (2019); Rogers et al. (2020); Madsen et al. (2022), these methods are not suitable for exploring fine-grained visual features.
@@
Next, we investigate memory retrieval, and uncover several key findings: (1) Common concepts, such as “Kangaroo”, emerge in early layers while less common ones, such as the animal “Dik-dik”, gradually emerge across the layers, with the most accurate representations predominantly occurring in the upper layers, as illustrated in Figure 2 (Right, top). (2) Fine details, like human facial features, materialize at later layers, as shown in Figure 2 (Right, bottom). (3) Knowledge retrieval is gradual, unfolding as computation progresses. This observation diverges from prior research on knowledge encoding in language models which characterizes knowledge as a localized attribute encoded in specific layers Geva et al. (2022); Meng et al. (2022); Arad et al. (2023). (4) Notably, there are discernible differences in memory retrieval patterns between the two text encoders: Deep Floyd’s T5 memory retrieval exhibits a more incremental behavior compared to Stable Diffusion’s CLIP. The disparities uncovered through our analyses suggest that factors such as architecture, pretraining objectives, or data may influence the encoding of knowledge or language representation within the models.
@@ 2 Diffusion Lens
Current text-to-images diffusion models comprise two main components Saharia et al. (2022); Ramesh et al. (2022): a language model used as a text encoder that takes the textual prompt as input and produces latent representations; and a diffusion model that is conditioned on the representations from the text encoder and generates an image from an initial input noise.
@@ 3 Experimental Setup
The experiments are performed on Stable Diffusion 2.1 (denoted SD, Rombach et al., 2022) and Deep Floyd (denoted DF, StabilityAI, 2023). SD is an open-source implementation of latent diffusion Rombach et al. (2022), with OpenCLIP-ViT/H Ilharco et al. (2021) as the text-encoder. DF is another open-source implementation of latent diffusion inspired by Saharia et al. (2022), with a frozen T5-XXL Raffel et al. (2020) as the text encoder. We usually only report the results on DF, unless there is a difference between the models, which we then discuss. The full results on SD are given in Appendix F.
@@ Results
Figure 3 shows the percentage of images that correctly generated the concepts for each category: an object, an object and a color, and two colored objects. Prompts describing a single object emerge the earliest, between layers 4 and 16, while prompts containing a color descriptors emerge in layers 16–20. Conjunction prompts emerge last, around layers 20–24. We observe a similar pattern for the preposition prompts, which we describe in Appendix A.1. As demonstrated in Figure 4, “A cow” is fully represented by layer 8, while “A yellow dolphin” does not correctly form until layer 16. Lastly, “A pink snail and an orange donut” only fully forms at much later layers, correctly matching the objects and colors at the final layer, 24.
@@
We continue with the complex prompts of two colored objects. Figure 5 aggregates the answers to illustrate the behavior of either or both objects appearing in intermediate layers. Colors often emerge first, with both colors often emerging in early layers in DF (in SD, the two objects appear before two colors). A single object is also gradually represented in layers 4-12. Notably, while the colors and one of the objects appear, the object is not necessarily generated in the correct color. This can be seen in the first example in Figure 6: While a raccoon and a rocket do appear, and the image contains both blue and pink elements, the rocket is not blue until the final layer. In some cases, we observe a mixture of concepts in early layers, as seen in the second example of Figure 6. Similarly, the bottom two examples in Figure 6 show prompts composing two objects and a proposition. As with colors, we observe that individual objects appear in early layers but the correct relation emerges much later. For example, “A cake on a cloud” generates images of both a cake and a cloud, with different relations; at layer 8 the cake is decorated with a cloud and in layer 20 the clouds are depicted as frosting. The correct relation is only generated at the final layer. These patterns suggest that the early layers of the text encoder behave like a “bag of concepts”, with a representation for each concept but no clear relations between them.
@@ Results.
Second, Table 1 presents information on the order of generation for both models, revealing that the sequence in which objects emerge during the computation process is determined by either their linear or their syntactic precedence, depending on the particular text encoder. In DF’s T5 text encoder, slightly over half of the instances feature the antecedent appearing at an earlier layer than the descendant, with a smaller fraction showing the opposite, and the rest indicating simultaneous appearances. This holds true regardless of linear order. Conversely, in SD’s Clip, the first noun tends to appear before the second more frequently, irrespective of the syntactic role.
@@ 5.1 Common and Uncommon Concepts
We investigate whether there is a difference in the generation process for prompts describing common and uncommon concepts, using a list of common and uncommon animals.333We use animal species as we found that models can correctly generate images of uncommon species, unlike uncommon objects and celebrities. Commonality in this context does not refer to the commonality of an animal in the world, but rather to its commonality in the training data. As a proxy to measure commonality in the training data, we utilized the average daily view statistics of Wikipedia pages from October 2022 to October 2023. An animal was deemed “common” if it had an average of 1500 visits per day on its Wikipedia page (e.g., kangaroo), while one having fewer than 800 visits per day was deemed “uncommon”. We verified this distinction by examining the frequencies of species names in the LAION2B-en dataset Schuhmann et al. (2022), extracted by Samuel et al. (2024), and found that the frequency of common species was greater than that of uncommon species with statistical significance (Appendix D).
@@ Results.
Though the plot for SD reveals a similar pattern (Appendix F), qualitative analysis reveals distinct knowledge retrieval patterns between the two models: In the case of DF’s T5, knowledge retrieval is gradual, unfolding as computation progresses (Figure 9). Layers generate animal, mammal, and ultimately construct a representation of the specific animal. However, SD’s text encoder, Clip, does not display a similar progression of retrieval. The model establishes the representation less gradually: The first layer with a meaningful image already closely resembles the final animal, with subsequent layers mainly refining its characteristics. These differences echo the syntactic findings in Section 4.2. They suggest that pretraining objectives, data, or model architecture might influence information organization, leading to distinct memory retrieval patterns.
@@ 6 Analyzing Model Failures
In this section, we delve into cases where the T2I model fails to generate images that align with the input prompt. We focus on prompts describing two entities with different colors, as these prompts lead to the highest failure rate in our experiments. Examples for failure cases are shown in figure 12. Examining the output images, the failures look similar: in both cases one or more entity was generated in the wrong color. However, using the Diffusion Lens, we reveal two different failure patterns: In the first example (“A green cat and a blue rocket”), both the blue rocket and the green cat are generated separately successfully in intermediate layers, while final output image fails to combine them into a single image. This suggests that the failure stems from an unsuccessful combination of the two concepts. In the second example (“A black bird and a green bear”), the bear consistently appears black across all intermediate layers, signifying that the model struggles to generate a green bear throughout the encoding process of the text. A possible explanation is that black bear is a type of animal, which might mean that the phrase “black bear” is common in the training data, thus the appearance of the phrase “black” in the prompt biases the model towards generating black bears. This analysis reveals two types of failures that occur in compound prompts: either (1) the model fails at coupling a particular concept and color because it is biased towards another color, or that (2) the model can successfully couple each concept and color but fails to combine them.
We inspect how often the two types of failures appear. Here, we examine only prompts that failed to generate a correct image for at least 75% of the generations from the final layer. For each entity, we count the number of times it appeared in the correct color in at least one early layer. We find that for 40% of the failure cases in DF (70% in SD), at least one of the entities did not appear at all in images from earlier layers (type 1). The remianing set of failures had the correct color for each of the objects appear at some point in the computation (type 2). An additional breakdown of failure cases per experiment can be found at Appendix C.
@@ 7 Related Work
A wide range of work has analyzed language model internals. We briefly mention a few directions and refer to existing surveys Belinkov and Glass (2019); Rogers et al. (2020); Madsen et al. (2022). Probing classifier are used to analyze whether internal representations correlate with external properties (e.g., Ettinger et al., 2016; Hupkes et al., 2018). However, probing has various inherent flaws such as memorization Belinkov (2022), and requires costly annotations for fine-grained analysis like the visual characteristics of a specific animal species or person. Interventions in representations measure how they impact a model’s prediction (e.g., Vig et al., 2020; Elazar et al., 2021; Meng et al., 2022), and while they offer powerful insights, they are also challenging to design Zhang and Nanda (2023) and limited in scope. In contrast, the Diffusion Lens proposes a simple yet generic mechanism to visually interpret intermediate representations without requiring additional data or training, enabling exploration of fine-grained visual features.
@@
Interpreting vision–language models.
Compared to unimodal models, research on interpretability in multimodal vision–language models is rather limited. Goh et al. (2021) found multi-modal neurons responding to specific concepts in CLIP Radford et al. (2021) and Gandelsman et al. (2023) decomposed CLIP’s image representations into text-based characteristics.
@@ A.1 Prepositions
We explore prepositions in prompts. We investigate how prompts, including certain relations, affect the generation process. These prompts are complex, challenging the compositional understanding of the T2I model. In particular, we examine the prepositions "on" and "in". Figure 13 illustrates the percentage of images that correctly generated the concepts for two categories: objects alone and objects with specified relation. Our findings reveal that prompts involving only one of the objects tend to perform well in the early layers of the model. However, more intricate prompts, including both objects and a relational context, emerge only in later layers of the model.
@@ Appendix B Annotation Process
Appendix C Model Failures
Figure 18 shows the percentage of failures for each experiment that had over 10 failures. We split failures to two types: complete failures when no layer generated a correct image through Diffusion Lens, and cases when at least one layer generated a correct image, but the top layer led to a failure (success then failure).
Generally, the percentage of failure cases (total height of each bar) is low, from 10% to 25% for most categories. Prompts about two colored objects have a higher failure rate. Importantly, in many failure cases, the representations in earlier layers lead to a correct generation via our method. Notably, in simple prompts (relations and colored objects), about 80% of the failures had successful generations at earlier layers. See Figure 19 for an example. Once more constraints are imposed (two colored objects), we have a lower rate of early success. Finally, for knowledge-related tasks (famous people, unfamiliar animals), there are very few cases of early success turned to failure. Presumably, when the model fails, it is mostly because it does not encode the information at all.
Appendix D Animals Experiment: Implementation Details
D.1 Animal classes used
@@
D.2 The full list of animals
@@
Appendix E Implementation Details
We implemented our code using Pytorch (Paszke et al., 2019) and Huggingface libraries (Wolf et al., 2020; von Platen et al., 2022). For each experiment, we generated four images (different seeds) for each layer, and we report the standard division over the seeds in all plots. We use Stable Diffusion v2-1 (CreativeML Open RAIL++-M License) (Rombach et al., 2022) and Deep Floyd (DeepFloyd-IF-License) StabilityAI (2023). We ran the experiments on the following GPUs: Nvidia A40, RTX 6000 Ada Generation, RTX A4000 and GeForce RTX 2080 Ti.
@@ Appendix E Implementation Details
E.1 Dependency parsing implementation
@@
Appendix F Results on Stable Diffusion

### v2

@@ 1 Introduction
The text-to-image (T2I) diffusion pipeline comprises two main elements: the text encoder and the diffusion model. The text encoder converts a textual prompt into a latent representation, while the diffusion model utilizes this representation to generate the corresponding image. Several recent studies have delved into the internal workings of the diffusion model and the cross-attention mechanism that connects the two components Tang et al. (2023); Hertz et al. (2023); Orgad et al. (2023); Chefer et al. (2023a). Yet, while the text encoder is a key component of the pipeline with a large effect on image quality and text-image alignment Saharia et al. (2022), its internal mechanisms remain unexplored. Moreover, while there is a wide range of work that has analyzed general language model internals Belinkov and Glass (2019); Rogers et al. (2020); Madsen et al. (2022), these methods are not suitable for exploring fine-grained visual features.
@@
Next, we investigate memory retrieval, and uncover several key findings: (1) Common concepts, such as “Kangaroo”, emerge in early layers while less common ones, such as the animal “Dik-dik”, gradually emerge across the layers, with the most accurate representations predominantly occurring in the upper layers, as illustrated in Figure 22 (Right, top). (2) Fine details, like human facial features, materialize at later layers, as shown in Figure 22, with the prompt “A photo of Steve Jobs”. (3) Knowledge retrieval is gradual, unfolding as computation progresses. This observation diverges from prior research on knowledge encoding in language models which characterizes knowledge as a localized attribute encoded in specific layers Geva et al. (2022); Meng et al. (2022); Arad et al. (2023). (4) Notably, there are discernible differences in memory retrieval patterns between the two text encoders: Deep Floyd’s T5 memory retrieval exhibits a more incremental behavior compared to Stable Diffusion’s CLIP. The disparities uncovered through our analyses suggest that factors such as architecture, pretraining objectives, or data may influence the encoding of knowledge or language representation within the models.
@@ 2 Diffusion Lens
Current text-to-image diffusion models comprise two main components Saharia et al. (2022); Ramesh et al. (2022): a language model used as a text encoder that takes the textual prompt as input and produces latent representations; and a diffusion model that is conditioned on the representations from the text encoder and generates an image from an initial input noise.
@@ 3 Experimental Setup
The experiments are performed on Stable Diffusion 2.1 (denoted SD, Rombach et al., 2022) and Deep Floyd (denoted DF, StabilityAI, 2023). SD is an open-source implementation of latent diffusion Rombach et al. (2022), with OpenCLIP-ViT/H Ilharco et al. (2021) as the text-encoder. DF is another open-source implementation of latent diffusion inspired by Saharia et al. (2022), with a frozen T5-XXL Raffel et al. (2020) as the text encoder. We usually only report the results on DF, unless there is a difference between the models, which we then discuss. The full results on SD are given in Appendix E.
@@ Results
Figure 3 shows the percentage of images that correctly generated the concepts for each category: an object, an object and a color, and two colored objects. Prompts describing a single object emerge the earliest, between layers 4 and 16, while prompts containing a color descriptors emerge in layers 16–20. Conjunction prompts emerge last, around layers 20–24. As demonstrated in Figure 4, “A cow” is fully represented by layer 8, while “A yellow dolphin” does not correctly form until layer 16. Lastly, “A pink snail and an orange donut” only fully forms at much later layers, correctly matching the objects and colors at the final layer, 24.
@@
We continue with the complex prompts of two colored objects. Figure 5 aggregates the answers to illustrate the behavior of either or both objects appearing in intermediate layers. Colors often emerge first, with both colors often emerging in early layers in DF (in SD, the two objects appear before two colors). A single object is also gradually represented in layers 4-12. Notably, while the colors and one of the objects appear, the object is not necessarily generated in the correct color. This can be seen in the first example in Figure 6: While a raccoon and a rocket do appear, and the image contains both blue and pink elements, the rocket is not blue until the final layer. In some cases, we observe a mixture of concepts in early layers, as seen in the second example of Figure 6. Similarly, the bottom two examples in Figure 6 show prompts composing two objects and a proposition. As with colors, we observe that individual objects appear in early layers but the correct relation emerges much later. For example, “A cake on a cloud” generates images of both a cake and a cloud, with different relations; at layer 8 the cake is decorated with a cloud and in layer 20 the clouds are depicted as frosting. The correct relation is only generated at the final layer. We provide a more detailed discussion of prepositions in Appendix A.1. The patterns we see in these prompts suggest that the early representations of the text encoder behave like a “bag of concepts”, with a representation for each concept but no clear relations between them.
@@ Results.
Second, Table 1 presents information on the order of generation for both models, revealing that the sequence in which objects emerge during the computation process is determined by either their linear or their syntactic precedence, depending on the particular text encoder. In DF’s T5 text encoder, slightly over half of the instances feature the antecedent appearing at an earlier layer than the descendant, with a smaller fraction showing the opposite, and the rest indicating simultaneous appearances. This holds true regardless of linear order. Conversely, in SD’s CLIP, the first noun tends to appear before the second more frequently, irrespective of the syntactic role. See Figure 7 for a qualitative example of this case.
@@ 5.1 Common and Uncommon Concepts
We investigate whether there is a difference in the generation process for prompts describing common and uncommon concepts, using a list of common and uncommon animals.333We use animal species as we found that models can correctly generate images of uncommon species, unlike uncommon objects and celebrities. Commonality in this context does not refer to the commonality of an animal in the world, but rather to its commonality in the training data. As a proxy to measure commonality in the training data, we utilized the average daily view statistics of Wikipedia pages from October 2022 to October 2023. An animal was deemed “common” if it had an average of 1500 visits per day on its Wikipedia page (e.g., kangaroo), while one having fewer than 800 visits per day was deemed “uncommon”. We verified this distinction by examining the frequencies of species names in the LAION2B-en dataset Schuhmann et al. (2022), extracted by Samuel et al. (2024), and found that the frequency of common species was greater than that of uncommon species with statistical significance (Appendix C).
@@ Results.
Though the plot for SD reveals a similar pattern (Appendix E), qualitative analysis reveals distinct knowledge retrieval patterns between the two models: In the case of DF’s T5, knowledge retrieval is gradual, unfolding as computation progresses (Figure 10). Layers generate animal, mammal, and ultimately construct a representation of the specific animal. However, SD’s text encoder, Clip, does not display a similar progression of retrieval. The model establishes the representation less gradually: The first layer with a meaningful image already closely resembles the final animal, with subsequent layers mainly refining its characteristics. These differences echo the syntactic findings in Section 4.2. They suggest that pretraining objectives, data, or model architecture might influence information organization, leading to distinct memory retrieval patterns.
@@ 6 Analyzing Model Failures
In this section, we delve into cases where the T2I model fails to generate images that align with the input prompt. First, we investigate all failure cases in all experiments. Figure 14 shows the percentage of failures for each experiment that had over 10 failures. We split failures to two types: complete failures where no layer generated a correct image through Diffusion Lens, and cases where at least one layer generated a correct image, but the top layer led to a failure (success then failure).
Generally, the percentage of failure cases (total height of each bar) is low, from 10% to 25% for most categories. Prompts about two colored objects have a higher failure rate. Importantly, in many failure cases, the representations in earlier layers lead to a correct generation via our method. Notably, in simple prompts (relations and colored objects), about 80% of the failures had successful generations at earlier layers – see Figure 15 for an example. Once more constraints are imposed (two colored objects), we have a lower rate of early success. Finally, for knowledge-related tasks (famous people, unfamiliar animals), there are very few cases of early success turned to failure. Presumably, when the model fails, it is mostly because it does not encode the information at all.
Next, we zoom in on prompts describing two entities with different colors, as these prompts lead to the highest failure rate in our experiments. Examples for failure cases are shown in figure 13. Examining the final layer output images, the failures look similar: in both cases one or more entity was generated in the wrong color. However, using the Diffusion Lens, we reveal two different failure patterns: In the first example (“A green cat and a blue rocket”), both the blue rocket and the green cat are generated separately successfully in intermediate layers, while final output image fails to combine them into a single image. This suggests that the failure stems from an unsuccessful combination of the two concepts. In the second example (“A black bird and a green bear”), the bear consistently appears black across all intermediate layers, signifying that the model struggles to generate a green bear throughout the encoding process of the text. A possible explanation is that black bear is a type of animal, which might mean that the phrase “black bear” is common in the training data, thus the appearance of the phrase “black” in the prompt biases the model towards generating black bears. This analysis reveals two different sources of failures that occur in compound prompts: either (1) the model fails at coupling a particular concept and color because it is biased towards another color, or that (2) the model can successfully couple each concept and color but fails to combine them.
We analyze how frequently each source of failure occurs, focusing on prompts that failed to generate a correct image from the final layer in at least 75% of cases. For each entity, we count the number of times it appeared in the correct color in at least one early layer. We find that for 40% of the failure cases in DF (70% in SD), at least one of the entities did not appear at all in images from earlier layers (type 1). The remaining set of failures had the correct color for each of the objects appear at some point in the computation (type 2).
@@ 7 Related Work
A wide range of work has analyzed language model internals. We briefly mention a few directions and refer to existing surveys Belinkov and Glass (2019); Rogers et al. (2020); Madsen et al. (2022); Ferrando et al. (2024). Probing classifier are used to analyze whether internal representations correlate with external properties (e.g., Ettinger et al., 2016; Hupkes et al., 2018). However, probing has various inherent flaws such as memorization Belinkov (2022), and requires costly annotations for fine-grained analysis like the visual characteristics of a specific animal species or person. Interventions in representations measure how they impact a model’s prediction (e.g., Vig et al., 2020; Elazar et al., 2021; Meng et al., 2022), and while they offer powerful insights, they are also challenging to design Zhang and Nanda (2023) and limited in scope. In contrast, the Diffusion Lens proposes a simple yet generic mechanism to visually interpret intermediate representations without requiring additional data or training, enabling exploration of fine-grained visual features.
@@
Interpreting vision-language models.
Compared to unimodal models, research on interpretability in multimodal vision-language models is rather limited. Goh et al. (2021) found multi-modal neurons responding to specific concepts in CLIP Radford et al. (2021) and Gandelsman et al. (2023) decomposed CLIP’s image representations into text-based characteristics.
@@ 8 Discussion and Conclusion
Acknowledgements
This research was supported by the Israel Science Foundation (grant 448/20), an Azrieli Foundation Early Career Faculty Fellowship, and an AI Alignment grant from Open Philanthropy. HO is supported by the Apple AIML PhD fellowship. DA is supported by the Ariane de Rothschild Women Doctoral Program.
@@ A.1 Prepositions
We explore prepositions in prompts. We investigate how prompts, including certain relations, affect the generation process. These prompts are complex, challenging the compositional understanding of the T2I model. In particular, we examine the prepositions "on" and "in". Figure 16 illustrates the percentage of images that correctly generated the concepts for tree categories: each of the objects alone and both objects with the correct relation between them. Our findings reveal that the emergence of each of the objects occurs at an early stage. However, both objects and their correct relation emerge only later in the text encoding.
@@ Appendix B Annotation Process
Appendix C Animals Experiment: Implementation Details
C.1 Animal classes used
@@
C.2 The full list of animals
@@
Appendix D Implementation Details
We implemented our code using Pytorch (Paszke et al., 2019) and Huggingface libraries (Wolf et al., 2020; von Platen et al., 2022). For each experiment, we generated four images (different seeds) for each layer, and we report the standard deviation over the seeds in all plots. We use Stable Diffusion v2-1 (CreativeML Open RAIL++-M License) (Rombach et al., 2022) and Deep Floyd (DeepFloyd-IF-License) StabilityAI (2023). We ran the experiments on the following GPUs: Nvidia A40, RTX 6000 Ada Generation, RTX A4000 and GeForce RTX 2080 Ti.
@@ Appendix D Implementation Details
D.1 Dependency parsing implementation
@@
Appendix E Results on Stable Diffusion